OpenStudy (goldeverything):

Translate this real-world problem into an equation: Jenna's dog Bruiser eats 5 times the number of ounces of food that her dog Princess eats. Bruiser also eats 4 milk-bones, for a total of nineteen ounces of food. How much food does Princess eat? five p plus four equals ninteen four p plus five equals ninteen nineteen plus four equals five p five p minus nineteen equals four

OpenStudy (anonymous):

4+5P=19 P=(19-4)/5=3 the answer is A

OpenStudy (anonymous):

THE ANSWER IS A

OpenStudy (anonymous):

4+5P=19 P=(19-4)/5=3 the answer is A